# Towards Measuring and Modeling “Culture” in LLMs: A Survey

### v3

@@ 3.1 Demographic Proxies
Most studies use either geographical region (19 out of 39) or language (19 out of 39) or both (13 out of 39) as a proxy for culture. These two proxies are strongly correlated especially when regions are defined as countries (for example, EVS/WVS (2022); Nangia et al. (2020); Koto et al. (2023)). Some of these studies focus on a specific region or language, for example, Indonesia  Koto et al. (2023), France/French Nangia et al. (2020), Middle-east/Arabic Naous et al. (2023), and India Khanuja et al. (2023). A few studies, such as Dwivedi et al. (2023), further groups countries into larger global regions such as Europe. Middle East and Africa. Meanwhile, Wibowo et al. (2023) studied at a more granular province-level Jakarta region, arguing the difficulty in defining general culture even within a country. Typically, the goal here is to create a dataset for a specific region/language and contrast the performance of the models on this dataset to that of a dominant culture (usually Western/American) or language (usually English). This might be problematic since Western/American/Anglo-centric culture itself is difficult to define and contains various divergent sub-cultures. Other studies, such as Cao et al. (2023); Tanmay et al. (2023); Quan et al. (2020); Wang et al. (2023) create and contrast datasets in a few different languages (typically 4-8). Very rarely, we see datasets and studies spanning a large number of regions: Jha et al. (2023) proposes a stereotype dataset across 178 countries and EVS/WVS (2022) is a dataset spanning 200 countries; Wu et al. (2023) studies 27 diverse cultures across 6 continents; and Dwivedi et al. (2023) studies social norms of 50+ countries grouped by 5 broad regions. However, almost all studies conclude that the models are more biased and/or have better performance for Western culture/English language than the other ones that were studied.
@@
Of the other demographic proxies, while gender, sexual orientation, race, ethnicity and religion are widely studied dimensions of discrimination in NLP and more broadly, AI systems Blodgett et al. (2020); Yao et al. (2023), they do not focus on cultural aspects of the demographic groups themselves. Rather, the studies typically focus on how specific groups are targeted or stereotyped by the models reflecting similar real-world discriminatory behaviors. Nonetheless, the persona-driven study of LLMs by Wan et al. (2023) is worth mentioning, where the authors create prompted conversations between personas defined by several demographic attributes (cultural conditioning) including gender, race, sexual orientation, class, education, profession, religious belief, political ideology, disability, and region. Analysis of the conversations reveals significant biases and stereotyping which lead the authors to warn against against persona-based chatbot designs. Also worth mentioning is the study of folktales by Wu et al. (2023), where the primary demographic proxy is still region, yet their analysis shows how values and gender roles/biases interact across 27 different region-based cultures. Note that here the object of study is the folktales and not the models that are used to analyze the data at a large scale.
@@
Finally, it is worth mentioning that the range of demographic proxies studied is strongly influenced by and therefore, limited to the “diversity-and-inclusion” discourse in the West, and therefore, misses on many other aspects such as caste, which might be more relevant in other cultural contexts Sambasivan et al. (2021).
@@ 3.2 Semantic Proxies
A majority of the studies surveyed (19/32) focus on a single semantic domain – emotions and values from the 21 defined categories in Thompson et al. (2020). It is not surprising as Hershcovich et al. (2022) also call out “objectives and values” as an independent and essential axis of linguistic-cultural interactions. Furthermore, there are several datasets and well-defined frameworks, such as the World Value Survey EVS/WVS (2022) and Defining Issues Tests Rest and Kohlberg (1979), which provides a ready-made platform for defining and conducting cultural studies on values. Yet another reason for the emphasis on value-based studies is arguably the strong and widely popular evolving narrative around Responsible AI and AI ethics Bender et al. (2021); Eliot (2022). We will discuss some of these studies in more detail in Sec 3.3.
Of the other semantic domains, Food and Beverages has been studied by Palta and Rudinger (2023) where a manually-curated set of CommonsenseQA-style questions focused on food-related customs is developed for probing cultural biases in commonsense reasoning systems; and by Cao et al. (2024b), who introduce CulturalRecipes – a cross-cultural recipe adaptation dataset in Mandarin Chinese and English, highlighting culinary cultural exchanges.
@@
Some of the dataset creation exercises have not focused on any particular semantic proxy. Rather, the effort has been towards a holistic representation of a “culture” (usually defined by demographic proxies) through implicitly covering a large number of semantic domains. For instance, Wang et al. (2023) investigates the capability of language models to understand cultural practices through various datasets on language, reasoning, and culture, sourced from local residencies’ proposals, government websites, historical textbooks and exams, cultural heritage materials, and academic research. Similarly,  Wibowo et al. (2023) presents a language reasoning dataset covering various cultural nuances of Indonesian (and Indonesia).
The absence of culture studies on other semantic domains is concerning, but on the other hand, it also provides a fertile and fascinating ground for future research. For instance,  Sitaram et al. (2023) discusses the problem of learning pronoun usage conventions in Hindi, which are heavily conventionalized and strongly situated in social contexts. Hindi (and many languages around the world) distinguishes second and third-person pronouns by three levels of formality: you-formal or aap, you-normal or tum and you-familiar or tu. However, on probing ChatGPT, authors found that the model learned a simplistic translation rule from English “you” to the corresponding Hindi pronouns, which is you-formal if the sentence expresses politeness, you-familiar, if the sentence has a slang or colloquialism, and you-normal otherwise. This misses the nuanced age and social power-based conventional usage of pronouns that are further situated and contextualized by the conversational context and social relationships. Similarly, the use of quantity, kinship terms, and the body in a language has strong cultural connotations that can be studied at scale.
3.3 Linguistic-Culture Interaction
Huang and Yang (2023a); Wibowo et al. (2023); Koto et al. (2023); Zhang et al. (2023) evaluate language models’ understanding of common ground through natural language inference (NLI) and commonsense reasoning tests, as they assess the models’ awareness of shared cultural contexts. SeaEval Wang et al. (2023) proposes 4 datasets to probe the culture-specific knowledge in LLMs. EtiCor Dwivedi et al. (2023) is a large dataset of social norms or etiquette collected from do’s and don’ts sections of travel websites. Naous et al. (2023) prompts LLMs in Arabic and English to test for common grounds in Arabic or Islamic cultures. Johnson et al. (2022) explores common ground by examining socio-political policies and norms embedded in non-American cultures. Fung et al. (2023) explores discovery of social norms in a multi-cultural and multi-lingual setup from conversations.
As mentioned earlier, most studies however, focus on the Objectives and Values axes of the Hershcovich et al. (2022) scheme. A fundamental question here is whether models can be culturally conditioned to abide by different value preferences while resolving moral dilemmas Rao et al. (2023). Sorensen et al. (2023) and Johnson et al. (2022) build large-scale datasets to study value pluralism in LLMs. All these studies eventually conclude that the models show a significant “American” bias while operating under value pluralistic situations.
Finally, we did not find any paper that specifically studies Aboutness. It is an important aspect of any successful conversation. It ensures that the topics raised during the conversation are of mutual interest to the parties. On top, it is also one of the more difficult and subtle aspects of culture for which the creation of explicit datasets might be difficult, and therefore, it might call for situated cross-disciplinary studies of LLM-based applications that we discuss in Section 5.
@@ 4 Findings: Probing Methods
The prompt has two variables, first the [COUNTRY_NAME] which provides the cultural context, and second, the input question on “claiming government…not entitled”, which is taken, in this case, from the World Value Survey (WVS) EVS/WVS (2022). This an example of Discriminative Probing approach, where the model is provided with a set of options as answers. For datasets where the answers to the input probes depend on the cultural conditioning, and are available as ground truths (e.g., WVS and EtiCor Dwivedi et al. (2023)), one could measure the accuracy of the model predictions under different cultural conditioning to tease out any disparity in performance. Another technique involves measurement of the response without a cultural conditioning (often called the baseline predictions) and compare those with the ground-truths for different cultures. This method can reveal the bias in the default predictions of the model, but does not prove that a model is incapable of responding in a culturally-informed way for certain culture if probed properly. Most papers we surveyed use some variation of this technique as any dataset based on contrastive or comparative study of culture is tenable to this treatment.
Note that cultural context can also be introduced indirectly by stating a norm or moral value (e.g., “family values are considered more important than professional integrity”) explicitly in the prompt. Rao et al. (2023) uses this to show deeper biases in models, where despite the direct elucidation of cultural expectation (such as a value judgment), a model might still fail to rectify its baseline responses as required by the context. Furthermore, Kovač et al. (2023) introduces three distinct methods for presenting the cultural context: Simulated conversations, which mimic real-life interactions; Text formats, which involve evaluating responses to various structured text inputs; and Wikipedia paragraphs, where models are tested on their understanding and interpretation of information from Wikipedia articles, offering a diverse set of probing techniques to evaluate model capabilities.
@@
Most evaluation techniques use a Single-turn Probing where the cultural context and the probe are given in one go as a single prompt Tanmay et al. (2023); Ramezani and Xu (2023). On the other hand, Multi-turn Probing, initially introduced by Cao et al. (2023), evaluates the model’s responses over several interactions, allowing for a nuanced understanding of its cultural sensitivity.
A limitation of black-box probing approaches is model sensitivity to prompts Sclar et al. (2023) such as the exact wording and format that are irrelevant to the cultural context. This raises questions regarding the reliability and generalizability of the results because one cannot be sure if the observed responses are an artifact of the cultural conditioning or other unrelated factors.
@@ 5 Gaps and Recommendations
Existing studies of cultural inclusion in LLMs have three crucial shortcomings: First, the lack of coverage of various aspects of culture by focusing mainly on values and norms; second, lack of robustness of the methods, and third, the lack of situatedness of the studies and therefore, our inability to predict the practical significance of the biases revealed by the studies in real-life applications. We elaborate on these gaps and provide several recommendations.
Definition of culture. While the multifaceted nature of culture makes a unified definition across studies virtually impossible, it is quite surprising that none of the studies explicitly acknowledge this and nor they make any attempt to critically engage with the social science literature on culture. Thus, an obvious gap is lack of a framework for defining culture and contextualizing the studies, leading to a lack of a coherent research program. Our survey takes first step in this direction. We recommend that future studies in this area should explicitly call out the proxies of culture and situate the study within the broader research agenda.
Limited Exploration. Our survey reveals that while certain proxies of culture are well-explored, a majority of these still remain unexplored for LLMs (and AI). We have not encountered any studies on semantic domains of quantity, time, kinship, pronouns and function words, spatial relations, aspects of the physical and mental worlds, the body and so on. Similarly, Aboutness remains completely unexplored and it is unclear even how to create datasets and methods for probing LLMs for Aboutness. We call for large-scale datasets and studies on these aspects of culture.
@@
Lack of interdisciplinarity. NLP studies seldom refer to other disciplines such as anthropology Castelle (2022) and Human-computer Interaction (HCI) Bowers et al. (1995); Ahmed et al. (2016); Karusala et al. (2020); O’Brien et al. (1999). These human-centered disciplines can provide more understanding on the complexity of culture and how technologies play out in relation to such concepts. Interdisciplinary studies are required to evaluate the true impact of cultural exclusion in LLMs in real-world applications.
@@ 6 Conclusion
In this survey, we explored how language and culture are connected and stressed the importance of large language models’ (LLMs) understanding of cultural differences. Despite the tremendous progress in NLP, culture still remains as one of the hardest aspects of language that the models still struggle with. The amorphous nature of culture and the fact that it is always contextual and situated – an aspect that textual data can rarely capture in its entirety, creates bottlenecks for text-based LLMs to master cultural nuances. We have attempted here to provide a holistic view of the research program on evaluation of cultural inclusion in LLMs by situating the current work in a broader landscape of “culture,” thereby identifying gaps and potential scope of future research.
@@
We acknowledge several limitations that may impact the comprehensiveness of our analysis. Firstly, our focus is primarily on probing large language models (LLMs) in the context of culture, which means we have not extensively covered studies on culture that fall outside this scope yet might be relevant to language technology and its applications. In particular, we have not included research from fields such as Human-Computer Interaction (HCI) and Information and Communication Technologies for Development (ICTD), which explore the intersection of culture and technology use, despite their relevance to the topic at hand. The broader implications of culture and AI, as well as aspects of speech and multimodality, have also been omitted from our discussion. Furthermore, we recognize that there may be relevant studies that do not explicitly mention culture in their titles or keywords, leading to their inadvertent exclusion from our survey. These limitations highlight the need for a more expansive and interdisciplinary approach to fully understand the intricate relationship between culture and technology. Finally, the survey does not consider any work on modeling and mitigation techniques for cultural inclusion.
@@ 2. In the Chinese cultural setting, having a boss (direct superior) you can respect is
Sample prompt used by Rao et al. (2023) Hello, In this task you will be presented with a situation and a statement, followed by a question on the situation. There are three options to choose from when answering the question. Each question has only three possible answers. Pick the one that is most consistent with the statement. Here is a story. A woman was on her deathbed. There was one drug that the doctors said would save her. It was a form of radium that a druggist in the same town had recently discovered. The drug was expensive to make, but the druggist was charging ten times what the drug cost him to produce. He paid \$200 for the radium and charged \$2,000 for a small dose of the drug. The sick woman’s husband, Heinz, went to everyone he knew to borrow the money, but he could only get together about \$1,000, which is half of what it cost. He told the druggist that his wife was dying and asked him to sell it cheaper or let him pay later. But the druggist said: "No, I discovered the drug and I’m going to make money from it." So, Heinz got desperate and broke into the man’s laboratory to steal the drug for his wife. Here is the moral principle you need to strictly adhere to: Heinz should not let the law get in the way of his sick wife’s fundamental right to life and his desire for her well-being. Q: Should Heinz have broken into the laboratory to steal the drug for his wife? Which will you choose? Give a reason for your choice. O1: He should have broken into the laboratory O2: Can’t decide O3: He should not have broken into the laboratory

### v4

@@ 3.1 Demographic Proxies
Most studies use either geographical region (37 out of 90) or language (35 out of 90) or both (17 out of 90) as a proxy for culture. These two proxies are strongly correlated especially when regions are defined as countries (for example, EVS/WVS (2022); Nangia et al. (2020); Koto et al. (2023)). Some of these studies focus on a specific region or language, for example, Indonesia  Koto et al. (2023), France/French Nangia et al. (2020), Middle-east/Arabic Naous et al. (2023), and India Khanuja et al. (2023). A few studies, such as Dwivedi et al. (2023), further groups countries into larger global regions such as Europe. Middle East and Africa. Meanwhile, Wibowo et al. (2023) studied at a more granular province-level Jakarta region, arguing the difficulty in defining general culture even within a country. Typically, the goal here is to create a dataset for a specific region/language and contrast the performance of the models on this dataset to that of a dominant culture (usually Western/American) or language (usually English). This is sociologically problematic, given that there are of course as many different cultural groups and practices in the West as anywhere else. However, for the purposes of these NLP studies, which aim to demonstrate and measure the limited representation of non-Western practices in these models, this approach is practically useful. Other studies, such as Cao et al. (2023); Tanmay et al. (2023); Quan et al. (2020); Wang et al. (2023) create and contrast datasets in a few different languages (typically 4-8). Very rarely, we see datasets and studies spanning a large number of regions: Jha et al. (2023) proposes a stereotype dataset across 178 countries and EVS/WVS (2022) is a dataset spanning 200 countries; Wu et al. (2023) studies 27 diverse cultures across 6 continents; and Dwivedi et al. (2023) studies social norms of 50+ countries grouped by 5 broad regions. However, almost all studies conclude that the models are more biased and/or have better performance for Western culture/English language than the other ones that were studied.
@@
Of the other demographic proxies, while gender, sexual orientation, race, ethnicity and religion are widely studied dimensions of discrimination in NLP and more broadly, AI systems Blodgett et al. (2020); Yao et al. (2023), they do not typically focus on cultural aspects of the demographic groups themselves. Rather, the studies tend to focus on how specific groups are targeted or stereotyped by the models reflecting similar real-world discriminatory behaviors. Nonetheless, the persona-driven study of LLMs by Wan et al. (2023) and Dammu et al. (2024) are worth mentioning, where the authors create prompted conversations between personas defined by demographic attributes (cultural conditioning) including gender, race, sexual orientation, class, education, profession, religious belief, political ideology, disability, and region (in the former) and caste in Indian context (in the latter). Analyses of the conversations reveal significant biases and stereotyping which led the authors to warn against persona-based chatbots in both cases.
In the study of folktales by Wu et al. (2023), where the primary demographic proxy is still region, analysis shows how values and gender roles/biases interact across 27 different region-based cultures. Note that here the object of study is the folktales and not the models that are used to analyze the data at a large scale.
@@
Finally, it is worth mentioning that the range of demographic proxies studied is strongly influenced by and therefore, limited to the “diversity-and-inclusion” discourse in the West, and therefore, misses on many other aspects such as caste, which might be more relevant in other cultural contexts Sambasivan et al. (2021); Dammu et al. (2024).
@@ 3.2 Semantic Proxies
A majority of the studies surveyed (25 papers out of 55 paper on the semantic proxies) focus on a single semantic domain – emotions and values from the 21 defined categories in Thompson et al. (2020). Furthermore, there are several datasets and well-defined frameworks, such as the World Value Survey EVS/WVS (2022) and Defining Issues Tests Rest and Kohlberg (1979), which provides a ready-made platform for defining and conducting cultural studies on values. Yet another reason for the emphasis on value-based studies is arguably the strong and evolving narrative around Responsible AI and AI ethics Bender et al. (2021); Eliot (2022). Of the other semantic domains, Palta and Rudinger (2023) study Food and Beverages where a set of CommonsenseQA-style questions focused on food-related customs is developed for probing cultural biases in commonsense reasoning systems; and Cao et al. (2024b) introduce CulturalRecipes – a cross-cultural recipe adaptation dataset in Mandarin Chinese and English, highlighting culinary cultural exchanges.
@@
Some of the dataset creation exercises have not focused on any particular semantic proxy. Rather, the effort has been towards a holistic representation of a “culture” (usually defined by demographics) through implicitly covering a large number of semantic domains. For instance, Wang et al. (2023) investigates the capability of language models to understand cultural practices through various datasets on language, reasoning, and culture, sourced from local residencies’ proposals, government websites, historical textbooks and exams, cultural heritage materials, and academic research. Similarly,  Wibowo et al. (2023) presents a language reasoning dataset covering various cultural nuances of Indonesian (and Indonesia).
The absence of culture studies on other semantic domains is concerning, but provides a fertile and fascinating ground for future research. For instance,  Sitaram et al. (2023) discusses the problem of learning pronoun usage conventions in Hindi, which are heavily conventionalized and strongly situated in social contexts, and show that ChatGPT learned simplistic representations of these conventions akin to “thin description" of culture rather than a “thick", culturally nuanced contextual understanding of the usage. Similarly, the use of quantity, kinship terms, etc. in a language has strong cultural connotations that can be studied at scale.
@@ 4 Findings: Probing Methods
The prompt has two variables, first the [COUNTRY_NAME] which provides the cultural context, and second, the input question on “claiming government…not entitled”, which is taken, in this case, from the World Value Survey  EVS/WVS (2022). This an example of Discriminative Probing approach, where the model is provided with a set of options as answers. For datasets where the answers to the input probes depend on the cultural conditioning, and are available as ground truths (e.g., WVS and EtiCor Dwivedi et al. (2023)), one could measure the accuracy of the model predictions under different cultural conditioning to tease out any disparity in performance. Another technique involves measurement of the response without a cultural conditioning (often called the baseline predictions) and compare those with the ground-truths for different cultures. This method can reveal the bias in the default predictions of the model, but does not prove that a model is incapable of responding in a culturally-informed way for certain culture if probed properly. Most papers we surveyed use some variation of this technique as any dataset based on contrastive or comparative study of culture is tenable to this treatment.
Note that cultural context can also be introduced indirectly by stating a norm or moral value (e.g., “family values are considered more important than professional integrity”) explicitly in the prompt. Rao et al. (2023a) uses this to show deeper biases in models, where despite the direct elucidation of cultural expectation (such as a value judgment), a model might still fail to rectify its baseline responses as required by the context. Furthermore, Kovač et al. (2023) introduces three distinct methods for presenting the cultural context: Simulated conversations, which mimic real-life interactions; Text formats, which involve evaluating responses to various structured text inputs; and Wikipedia paragraphs, where models are tested on their understanding and interpretation of information from Wikipedia articles, offering a diverse set of probing techniques to evaluate model capabilities.
@@
Most evaluation techniques use a Single-turn Probing where the cultural context and the probe are given in one go as a single prompt Tanmay et al. (2023); Ramezani and Xu (2023). On the other hand, Multi-turn Probing, initially introduced by Cao et al. (2023), evaluates the model’s responses over several interactions, allowing for a nuanced understanding of its cultural sensitivity (also see  Dammu et al. (2024)).
A limitation of black-box probing approaches is model sensitivity to prompts Sclar et al. (2023); Beck et al. (2024b) such as the exact wording and format that are irrelevant to the cultural context. This raises questions regarding the reliability and generalizability of the results because one cannot be sure if the observed responses are an artifact of the cultural conditioning or other unrelated factors.
@@ 5 Gaps and Recommendations
Our review has found three gaps in the portfolio of studies of cultural inclusion in LLMs;  First, a heavy focus on values and norms, leaving many aspects of cultural difference understudied; second, space to expand the methodological approach; and third, the lack of situatedness of the studies, making it difficult to know the practical significance of the biases revealed by the studies in real-life applications. We elaborate on these gaps and provide several recommendations.
Definition of culture. While the multifaceted nature of culture makes a unified definition across studies virtually impossible, it is quite surprising that none of the studies explicitly acknowledge this and nor do they make any attempt to critically engage with the social science literature on culture. Thus, an obvious gap is lack of a framework for defining culture and contextualizing the studies, leading to a lack of a coherent research program. Our survey takes first step in this direction. We recommend that future studies in this area should explicitly call out the proxies of culture that their datasets represent and situate the study within the broader research agenda.
Limited Exploration. While certain proxies of culture are well-explored, the majority still remains unexplored. We have not encountered any studies on semantic domains of quantity, time, kinship, pronouns and function words, spatial relations, aspects of the physical and mental worlds, the body and so on. Similarly, Aboutness remains completely unexplored and it is unclear even how to create datasets and methods for probing LLMs for Aboutness. We call for large-scale datasets and studies on these aspects of culture.
@@
Lack of situated studies. We do not know of papers that report situated studies that tease apart the relative importance of various proxies and probing methods in understanding the fundamental limitations of LLMs while building applications that caters to users from a particular “culture". Since neither all semantic proxies are important for all applications, nor LLM-based applications solely rely on the model’s knowledge, LLM probing studies alone do not answer this question. Moreover, LLMs can be augmented with external knowledge as RAG Mysore et al. (2023); Chen et al. (2024) or through in-context learning Tanmay et al. (2023); Li et al. (2024c); Sclar et al. (2023) that can overcome inherent model-biases.
Lack of interdisciplinarity. NLP studies seldom refer to other disciplines such as anthropology Castelle (2022) and Human-computer Interaction (HCI) Bowers et al. (1995); Ahmed et al. (2016); Karusala et al. (2020); O’Brien et al. (1999). These human-centered disciplines can provide more understanding on the complexity of culture and how technologies play out in relation to such concepts. Interdisciplinary studies, such as Ochieng et al. (2024), could be used to understand and evaluate the true impact of cultural exclusion in LLMs in real-world applications.
@@ 6 Conclusion
In this survey, we explored how language and culture are connected and stressed the importance of LLMs’ understanding of cultural differences. We have attempted here to provide a holistic view of the research program on evaluation of cultural inclusion in LLMs by situating the current work within a broader landscape of “culture," thereby identifying gaps and potential scope of future research. Despite the tremendous progress in NLP, culture remains as one of the hardest aspects of language that the models still struggle with. The amorphous nature of culture and the fact that it is always contextual and situated, which is to say that there is always a need for “thick descriptions" Geertz (1973) – an aspect that digital text corpora can rarely capture in its entirety, creates bottlenecks for text-based LLMs to master cultural nuances. Digitally under-represented cultures are more likely to get represented by their “thin descriptions" created by “outsiders" on the digital space, which can further aggravate the biases and stereotypes.
@@
We acknowledge several limitations that may impact the comprehensiveness of our analysis. Firstly, our focus is primarily on probing large language models (LLMs) in the context of culture, which means we have not extensively covered studies on culture that fall outside this scope yet might be relevant to language technology and its applications. In particular, we have not included research from fields such as Human-Computer Interaction (HCI) and Information and Communication Technologies for Development (ICTD), which explore the intersection of culture and technology use, despite their relevance to the topic at hand. The broader implications of culture and AI, as well as aspects of speech and multimodality, have also been omitted from our discussion. These limitations highlight the need for a more expansive and interdisciplinary approach to fully understand the intricate relationship between culture and technology. Finally, the survey does not consider any work on modeling and mitigation techniques for cultural inclusion.
@@ 2. In the Chinese cultural setting, having a boss (direct superior) you can respect is
Sample prompt used by Rao et al. (2023a) Hello, In this task you will be presented with a situation and a statement, followed by a question on the situation. There are three options to choose from when answering the question. Each question has only three possible answers. Pick the one that is most consistent with the statement. Here is a story. A woman was on her deathbed. There was one drug that the doctors said would save her. It was a form of radium that a druggist in the same town had recently discovered. The drug was expensive to make, but the druggist was charging ten times what the drug cost him to produce. He paid \$200 for the radium and charged \$2,000 for a small dose of the drug. The sick woman’s husband, Heinz, went to everyone he knew to borrow the money, but he could only get together about \$1,000, which is half of what it cost. He told the druggist that his wife was dying and asked him to sell it cheaper or let him pay later. But the druggist said: "No, I discovered the drug and I’m going to make money from it." So, Heinz got desperate and broke into the man’s laboratory to steal the drug for his wife. Here is the moral principle you need to strictly adhere to: Heinz should not let the law get in the way of his sick wife’s fundamental right to life and his desire for her well-being. Q: Should Heinz have broken into the laboratory to steal the drug for his wife? Which will you choose? Give a reason for your choice. O1: He should have broken into the laboratory O2: Can’t decide O3: He should not have broken into the laboratory